OpenStudy (josh0404):

help me to solve this, find the real solution 36x^(-(4)/(3))-13x(-2)/(3)+1=0

OpenStudy (josh0404):

@518nad help me

OpenStudy (josh0404):

@jim_thompson5910

OpenStudy (sshayer):

\[36 x ^{\frac{ -4 }{ 3 }}-13 x ^{\frac{ -2 }{ 3 }}+1=0\] Am I correct?

OpenStudy (sshayer):

If i am correct ,then put \[x ^{\frac{ -2 }{ 3 }}=t\] squaring \[\left( x ^{\frac{ -2 }{ 3 }} \right)^2=t^2,x ^{\frac{ -4 }{ 3 }}=t^2\] \[36 t^2-13t+1=0\] \[36 t^2-9 t-4 t+1=0\] \[9 t(4 t-1)-1(4t-1)=0\] \[(4t-1)(9t-1)=0,t=\frac{ 1 }{ 4 },\frac{ 1 }{ 9 }\] \[t=\frac{ 1 }{ 4 },x ^{\frac{ -2 }{ 3 }}=\frac{ 1 }{ 4 }\] \[\frac{ 1 }{ x ^{\frac{ 2 }{ 3 }} }=\frac{ 1 }{ 4 },\] \[x ^{\frac{ 2 }{ 3 }}=4 \] cubing both sides \[\left( x ^{\frac{ 2 }{ 3 }} \right)^3=4^3=64,x^2=64,x=\pm 8\] similarly when \[t=\frac{ 1 }{ 9 },x ^{\frac{ -2 }{ 3 }}=\frac{ 1 }{ 9 },\frac{ 1 }{ x ^{\frac{ 2 }{ 3 }} }=\frac{ 1 }{ 9 }\] \[x ^{\frac{ 2 }{ 3 }}=9,cubing\] \[x^2=9^3=729,x= \pm 27\]

OpenStudy (josh0404):

thank you so much

OpenStudy (josh0404):

@sshayer help me please

OpenStudy (josh0404):

help on this one i know, you know how to solve this one. 1)(1)/(1)+2(x+1)(1)/(5)=2

OpenStudy (josh0404):

3(x+1)(1)/(1)+2(x+1)(1)/(5)=2

OpenStudy (josh0404):

help me please

OpenStudy (sshayer):

your statement is ambiguous is it \[\frac{ 3\left( x+1 \right) }{ 1+2\left( x+1 \right)\frac{ 1 }{ 5 } }?\]

OpenStudy (josh0404):

\[3(x+1)\frac{ 1 }{ 1 }+2(x+1)\frac{ 1 }{ 5 }=2\]

OpenStudy (josh0404):

help on this one please?

OpenStudy (josh0404):

@brooke..help00

OpenStudy (sshayer):

\[\left( x+1 \right)\left( 3+\frac{ 2 }{ 5 } \right)=2\] \[\left( x+1 \right)\left( \frac{ 15+2 }{ 5 } \right)=2\] \[x+1=2 \times \frac{ 5 }{ 17 }=\frac{ 10 }{ 17 }\] \[x=\frac{ 10 }{ 17 }-1=?\]